OpenStudy (anonymous):

Find the distance between the following pairs of points - (a cos A, a sin A) and (a cos B, a sin B)

OpenStudy (anonymous):

pythagoras for this one (aka "distance formula") \[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

Right, but what do you get after substitution & simplification?

OpenStudy (phi):

how about you post what you get after substitution ? (as a first step)

OpenStudy (anonymous):

What I got is.....\[a.\sqrt{2(1-\cos(A-B))}\] but the answer key shows..... 2a.sin((A-B)/2)

OpenStudy (phi):

your answer is correct. But they used some trig identities to change it. Do you know the half-angle formulas? http://www.intmath.com/analytic-trigonometry/4-half-angle-formulas.php often written as \[ \sin\left(\frac{x}{2}\right) = \sqrt{\frac{1-\cos(x)}{2}}\] if we square both sides, we have \[ \sin^2\left(\frac{x}{2}\right) = \frac{1-\cos(x)}{2}\\ 2\sin^2\left(\frac{x}{2}\right) = 1-\cos(x) \] if we multiply both sides by 2 we get \[ 4\sin^2\left(\frac{x}{2}\right) = 2(1-\cos(x)) \] now take the square root of both sides \[ 2 \sin\left(\frac{x}{2}\right) = \sqrt{2(1-\cos(x))}\] use this result with x = A-B

OpenStudy (anonymous):

Thank you Phi! shall brush up half angle formula...